OpenStudy (anonymous):

if a:b=2:3 and b:c=5:3, then (a+b+c)/(a-b+c)=? @terenzreignz

terenzreignz (terenzreignz):

How do you go about this? \[\Large \frac{a}b= \frac23\]\[\Large \frac{c}b=\frac35\] This might help ^

terenzreignz (terenzreignz):

Still no clue?

terenzreignz (terenzreignz):

\[\huge \frac{a+b+c}{a-b+c}\cdot \frac{\frac1b}{\frac1b}\] where, of course,\[\huge \frac{\frac1b}{\frac1b}=1\]

terenzreignz (terenzreignz):

once again, @kryton1212 I need you to make your presence felt :P Could you please simplify this? \[\huge \frac{a+b+c}{a-b+c}\cdot \frac{\frac1b}{\frac1b}\]

OpenStudy (anonymous):

(a+b+c)/(a-b+c) ??? lol

terenzreignz (terenzreignz):

Well of course... -.- But I need you to distribute the 1/b over both the numerator and the denominator...

OpenStudy (anonymous):

what do you mean?

terenzreignz (terenzreignz):

@kryton1212 ? Stay with me here... distribute the 1/b as instructed above ^

OpenStudy (anonymous):

i cannot get it...........

terenzreignz (terenzreignz):

I'll do the numerator, you do the denominator... \[\huge \frac{a+b+c}{a-b+c}\cdot \frac{\frac1b}{\frac1b}= \frac{\frac{a}b+\frac{b}b+\frac{c}b}{\color{red}?}\] Do you get it now?

OpenStudy (anonymous):

lol ?=(a/b)-(b/b)+(c/b)

terenzreignz (terenzreignz):

Yes. \[\huge \frac{\frac{a}b+\frac{b}b+\frac{c}b}{\frac{a}b-\frac{b}b+\frac{c}b}\] Now, clearly, \[\Large \frac{b}b=1\] What about \[\Large \frac{a}b\] and \[\Large \frac{c}b\]?

OpenStudy (anonymous):

(a+c)/b

terenzreignz (terenzreignz):

no... I meant, what is a/b = ? and what is c/b = ?

OpenStudy (anonymous):

a/b=2/3 c/b=3/5

terenzreignz (terenzreignz):

Yes... now \[\huge \frac{\frac{a}b+\frac{b}b+\frac{c}b}{\frac{a}b-\frac{b}b+\frac{c}b}= \frac{\frac{2}3+1+ \frac35}{\frac23-1+\frac35}\] I trust you can simplify it from here?

OpenStudy (anonymous):

17/2

OpenStudy (anonymous):

ohh, thank you very much!

terenzreignz (terenzreignz):

no problem.